OpenStudy (anonymous):

Simplify the expression: sqrt of -25/(5-2i)+(1-4i)

OpenStudy (anonymous):

\[\sqrt{-25}\div(5-2i)+(1-4i)\] for clarification

OpenStudy (anonymous):

\[\frac{ \sqrt{-25} }{ (5-2i)+(1-4i) }\] OR \[\frac{ \sqrt{-25} }{ (5-2i) }(1-4i)\]

OpenStudy (anonymous):

Whoops, forgot a plus sign in the second one.

OpenStudy (anonymous):

Ok so would I multiply the numerator and denominator by 1-4i?

OpenStudy (anonymous):

No no, sorry I was asking to clarify how to problem is written. :/

OpenStudy (anonymous):

\[\frac{ \sqrt{-25} }{ 5-2i }+(1-4i)\] That is what i meant to write.

OpenStudy (anonymous):

Is it that one, or the first one i wrote?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

Okay, that makes things a bit easier. The parentheses are meaningless here. You can easily simplify the denominator. We can also simplify the numerator. What do you think those would be?

OpenStudy (anonymous):

5i is the numerator and 6-6i is the numerator?

OpenStudy (anonymous):

sounds good to me.

OpenStudy (anonymous):

Now I'm not sure how much more "simple" you need to take it. Usually you try to get rid of the complex number from the denominator like what you usually do if you have a radical in the denominator.

OpenStudy (anonymous):

so do i multiply the numerator and denominator by the the conjugate of the denomiator?

OpenStudy (anonymous):

To do this, you need to multiply it by it's complex conjugate. Have you heard about that before?

OpenStudy (anonymous):

HAHA, that's great. We were on the same page and typing the same thing. Yes, that's exactly what you'd do.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so how would I do that exaclty?

OpenStudy (anonymous):

\[\frac{ 5i }{ 6-6i }*\frac{ 6+6i }{ 6+6i }\] The \[\frac{ 6+6i }{ 6+6i }\] is how we multiply by the complex conjugate. You'll notice it's just a "fancy" way of multiplying by one. This is probably one of the most important "tricks" you do in math, along with adding "zero" in fancy ways.

OpenStudy (anonymous):

What would the answer be? 30i+30i^2/72?

OpenStudy (anonymous):

This is where i get lost?

OpenStudy (anonymous):

which can be simplified. :D Do you know what i^2 is?

OpenStudy (anonymous):

also there are some common factors on top and bottom which can reduce everything.

OpenStudy (anonymous):

i^2=-1

OpenStudy (anonymous):

okay so what would my answer be?

OpenStudy (anonymous):

I don't want to just give you the answer. You're 100% correct so far though. Lets put everything you have together though. \[\frac{ 30i+(-30) }{ 72 } = \frac{ 30i-30 }{ 72 }\] Now lets start with 2. Does 2 go into 30 and 72?

OpenStudy (anonymous):

yes but wouldn't 6 be a better choice?

OpenStudy (anonymous):

It could be, but I always start with the lowest term possible. 2 is always a safe bet and if 2 doesn't work, try 3....and so on.

OpenStudy (anonymous):

is -5+5i/12 my answer?

OpenStudy (anonymous):

Yup!